Lee purchased a stock one year ago for $24. The stock is now worth $33, and the total return to Lee for owning the stock was 0.38. What is the dollar amount of dividends that he received for owning the stock during the year? Round to two decimal places.
Answer #1

Total return=(End value-Beginning value+Dividend)/Beginning value  

0.38=(33-24+Dividend)/24

(0.38*24)=9+Dividend

9.12=9+Dividend

Dividend=9.12-9

=$0.12